I was given a list of chemicals and was told I needed to explain the process. I know this is a disappearing ink problem and its fairly straight forward but I'm lost on some parts. I have solid NaHSO4 and needed to create a 5mL of 3M NaHSO4 solution. How do I do this? I have solid Na2S2O5 and need to make 10mL of 0.5M Na2S2O5 solution. How do I do this?
Answer #1

Now, Number of Moles of NaHSO4 needed = 3*(5/1000) = 15*10-3

Mass of NaHSO4 needed = 15*10-3 * 120 = 1.8 g

So take 5 mL water and add 1.8g NaHSO4 to it and you got your required 5 mL 3M NaHSO4 Solution

Moles of Na2S2O5 needed = 0.5*(10/1000) = 5*10-3

Mass of Na2S2O5 needed = 5*10-3 * 190 = 0.95 g

So take 10 mL water and add 0.95 g Na2S2O5 to it to get your required 10 mL 0.5M Na2S2O5 solution.